# Towards Realistic Data Generation for Real-World Super-Resolution

### v3

@@ 1 Introduction
Real-world image Super-Resolution (Real SR) is a fundamental problem in image processing, aiming to enhance the resolution and quality of images in real-world scenarios [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. It has a wide range of applications across various fields, including photography [18] and medical imaging [23], which enhance human visual perception and the robustness of vision systems [24, 25, 26, 27]. However, traditional bicubic-interpolation-based Real SR methods have proven less effective in complex real-world scenarios due to the significant discrepancy between the bicubic pattern and real degradation [1, 3, 28, 29, 30]. Consequently, substantial efforts have been directed towards developing methods for generating more realistic data to improve the generalization ability of Real SR models [28, 31, 32, 33, 34, 35, 36, 37, 38].
@@
To explore what kind of data contributes most to the SR model’s capability, we synthesize different sets of training data using [39] and evaluate the performance of FSRCNN [40]. As shown in the top left of Figure 1, the red triangles represent test samples of the target domain, while rectangular boxes indicate training sets with different blur kernels and noise levels. The results clearly demonstrate that the closer the training data distribution is to the test data, the better the model’s performance. This underscores the importance of designing a method that can adaptively generate accurate data for different target domains. Therefore, an ideal data generation system for Real SR should meet the following criteria: I) Large-scale, to satisfy the extensive data requirements for training deep learning models [41, 42, 43, 44]; II) Realistic, to enable Real SR models to accurately learn the characteristics of real-world degradation  [28, 31, 32]; and III) Adaptive, to flexibly generate data with arbitrary given degradation patterns, improving generalization in target domains [3, 34].
@@
Existing data generation methods for Real SR [28, 31, 32, 33, 42, 45, 46, 47, 48, 49], as shown in the bottom left of Figure 1, can be broadly categorized into the following: a) Manual Collection via Focal Length Adjustment: This approach involves using digital single-lens reflex cameras (DSLRs) with varying focal lengths to capture images, followed by alignment [28, 31]. While it can produce realistic paired data, it is labor-intensive and often results in scene monotony and image misalignment, failing to meet the large-scale data requirement. b) Hand-Crafted Physical-Based Degradation Modeling: This method employs various degradation models (e.g., noise, blur, bicubic, JPEG) applied in single-order or higher-order combinations [32, 33]. Although efficient in generating large data quantities, the synthetic data often fails to accurately reflect the complex degradation patterns of real-world images, and its lack of adaptability to specific domains limits effectiveness. c) Learning-Based Methods: Techniques involving Generative Adversarial Networks (GANs) [35, 45] and diffusion models [46] are proposed to simulate realistic real-world degradation for low-resolution (LR) images. While these methods produce more realistic data compared to hand-crafted approaches, they often struggle to generalize to new and diverse domains, limiting their applicability in real-world scenarios. In summary, existing data generation methods face challenges in achieving both realism and adaptability: tailoring models to specific target domains may hinder their adaptability to new domains, and vice versa. Overcoming these challenges is crucial for advancing the field of Real SR.
@@
In this paper, we introduce the unsupervised learning framework Realistic Decoupled Data Generator (RealDGen) to meet the large-scale, realistic, and adaptive data generation criteria. RealDGen enhances degradation realism and content fidelity by separately modeling content and degradation through unsupervised learning and integrating them into a diffusion model to generate paired data. The training involves two steps: first, pre-training degradation and content extractors using contrastive and reconstruction learning strategies to improve representation robustness; second, using these pre-trained extractors to condition the diffusion model with real LR degradation and content representations. To improve generalization to unknown LR distributions, the partial parameters of the extractors are fine-tuned. During data generation, unpaired HR and real LR images are used to extract content and degradation representations, which are then combined in the diffusion model. This process produces data that marries HR content with arbitrary LR degradation, improving adaptability to new domains. Extensive experiments show that RealDGen outperforms previous methods in generating realistic paired data and enhancing the performance of SR models in real-world scenarios.
@@
Figure 1: The top left and bottom left are SR performance on different train data and degradation distribution of different methods. The right is our unsupervised data generation framework RealDGen.
@@
Figure 2: An overview of our unsupervised training pipeline. We first train on the content and degradation extractors, then train Decoupled DDPM while fine-tuning the partial parameters of the extractors. RealDGen adaptively generates realistic LR images with arbitrary given real LR and unpaired HR conditions.
@@ 2.1 Content and Degradation Extractor Pre-training
To capture content and degradation representations, we propose dedicated degradation and content extractors, denoted as $E_{deg}$ and $E_{cont}$, respectively. We employ reconstruction learning for training $E_{cont}$, as shown in Figure 2 (a). Specifically, for a given high-resolution (HR) image $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, we degrade it to a low-resolution (LR) counterpart $\mathcal{X}_{lr}\in\mathbb{R}^{C\times h\times w}$ by Real-RSRGAN [33] degradation model $\mathcal{D}$ with diverse synthetic degradations. Subsequently, $E_{cont}$ is engaged to extract the content representation $F_{{cont}}$ from $\mathcal{X}_{lr}$. Thereafter, a HR reconstruction network $Rec_{hr}$ is harnessed to reconstruct the HR image $\hat{\mathcal{X}}\in\mathbb{R}^{C\times H\times W}$ from $F_{{cont}}$, as follows:
@@
$\mathcal{L}_{{rh}}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\mathcal{X}}_{i}-\mathcal{X}_{i}\right)^{2}\,,$ (2)
where $N$ denotes the batch size, which we set to 64. After training, $E_{cont}$ learns a robust content representation under diverse synthetic degradations.
Considering the variability of degradation in diverse scenarios and imaging devices in the real world, we advocate for contrastive learning to curate positive and negative samples for training $E_{deg}$, which guarantees the uniqueness of the degradation representations. Specifically, for a HR image $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, we generate an LR image $\mathcal{X}_{{lr}}\in\mathbb{R}^{C\times h\times w}$ by $\mathcal{D}$ with parameter $\theta$ as the anchor $\mathcal{X}_{{anc}}$. We further obtain a set of negative samples $\mathcal{X}_{{neg}_{i}}\in\mathbb{R}^{C\times h\times w}$ by applying $\mathcal{D}$ with different parameters $\theta_{i}$ to $\mathcal{X}$, and a set of positive samples $\mathcal{X}_{{pos}_{i}}\in\mathbb{R}^{C\times h\times w}$ by applying the degradation $\mathcal{D}$ with same parameter $\theta$ to different HR images $\mathcal{X}^{\prime}_{i}\in\mathbb{R}^{C\times H\times W}$, as follows:
$\begin{array}[]{c}\mathcal{X}_{{neg}}=\{\mathcal{D}\left(\mathcal{X},\theta_{1}\right),\mathcal{D}\left(\mathcal{X},\theta_{2}\right),\ldots,\mathcal{D}\left(\mathcal{X},\theta_{n}\right)\}\\ \mathcal{X}_{{pos}}=\{\mathcal{D}\left(\mathcal{X}^{\prime}_{1},\theta\right),\mathcal{D}\left(\mathcal{X}^{\prime}_{2},\theta\right),\ldots,\mathcal{D}\left(\mathcal{X}^{\prime}_{n},\theta\right))\}\,.\end{array}$ (3)
@@
$\small\mathcal{L}_{{cl}}=\sum_{i=1}^{n}\max\left(0,d\left(E_{deg}\left(\mathcal{X}_{{anc}}\right),E_{deg}\left(\mathcal{X}_{{pos}_{i}}\right)\right)-d\left(E_{deg}\left(\mathcal{X}_{{anc}}\right),E_{deg}\left(\mathcal{X}_{{neg}_{i}}\right)\right)+{margin}\right)\,,$ (4)
where $d$ symbolizes the L2 distance, $n$ is the number of samples, and we empirically set $n$ and $margin$ to $3$ and $0.01$, respectively. Furthermore, to drive $E_{deg}$ to learn the complete degradation representations, we utilize the reconstruction learning strategy as aforementioned for supervising $E_{deg}$. Specifically, we utilize the pre-trained $E_{cont}$ to learn the content representation of $\mathcal{X}_{{hr}}$ and $E_{deg}$ to learn the degradation representations of $\mathcal{X}_{{lr}}$. We then employ a low-resolution reconstruction network, $Rec_{lr}$, to combine these representations and reconstruct the LR image, $\hat{\mathcal{X}}_{{lr}}$, as follows:
@@
After training with well-designed learning strategies, the proposed encoders effectively capture robust content and degradation representations. More details are provided in Section A.1 and Section A.2 of the appendix.
@@
5:  $\mathbf{F_{cont}}=E_{cont}\left(\mathbf{x}_{lr}\right)$
6:  $\mathbf{F_{deg}}=E_{deg}\left(\mathbf{x}_{lr}\right)$
7:  $\mathbf{c}=\mathcal{M}\left(\mathbf{F_{cont}},\mathbf{F_{deg}}\right)$
@@ 2.2 Decoupled DDPM Training and Extractor Fine-tuning
We introduce a content-degradation Decoupled Diffusion Probabilistic Model (Decoupled DDPM) to generate real LR images. In detail, given real LR images ${\mathcal{X}}_{{lr}}$ from real-world $q$, we extract their content representation $F_{cont}$ and degradation representation $F_{deg}$ by pre-trained $E_{cont}$ and $E_{deg}$, respectively. To enhance the generalization of $E_{deg}$ and $E_{cont}$ on unseen real distributions, partial parameters are fine-tuned, as shown in Figure 2. A modulation block $\mathcal{M}$ is introduced to adequately incorporate degradation representation into the content, formulated as:
$\mathbf{c}=\mathcal{M}\left(E_{deg}\left(\mathbf{x}_{lr}\right),E_{cont}\left(\mathbf{x}_{hr}\right)\right)\,.$ (7)
Then, this fused image representation is utilized as a condition $\mathbf{c}$ for training our Decouple DDPM to generate LR images. To make it clear, we illustrate the detailed training procedure of the Decouple DDPM, as shown in Algorithm 1. A detailed presentation of $\mathcal{M}$ and fine-tuning is provided in Sections A.3A.4 and  A.5 of the appendix.
@@ 2.3 Data Generation
We propose a novel strategy to generate realistic LR images using unpaired LR and HR images by decoupling content and degradation. First, we extract the degradation representation from a real-world LR image and the content representation from an HR image. These representations are combined in the modulation module $\mathcal{M}$ to serve as the condition $\mathbf{c}$ for the diffusion model to generate LR images. The generated LR images retain the content of the HR image and the degradation of the real LR image, as shown on the right of Figure 1. To enhance fidelity, following [50], we denoise from an initial LR image $\mathbf{x}_{t}$ with $t$ steps of noise rather than from pure Gaussian noise. This initial LR image is degraded by ${\mathcal{D}}$. In Sec.3.4, we analyze the step number $T$. Details of our data generation pipeline are in Algorithm 2.
Although using content and degradation conditions improves the controllability and fidelity of Decoupled DDPM, the inherent stochasticity of the diffusion model [51, 52] can still introduce tiny artifacts and content distortion. To mitigate this, we propose a filtering mechanism. For each generated LR image, we re-extract content and degradation representations, then calculate the degradation error with the real LR image and the content error with the HR image. By selecting samples with the smallest errors, we reduce diffusion stochasticity and produce higher-fidelity LR images. More details will be presented in Section A.6 of the appendix.
@@ 3.1 Experiments Settings
Training Details. We have collected about 152,000 real low-resolution images from both public datasets [31, 28, 49] and those captured using smartphones. We train our extractors and Decouple DDPM on 16 NVIDIA V100 GPUs and set the learning rate and batch size to $1\times 10^{-4}$ and $16$, respectively. More details of the training setting are presented in Section  A.1 and  A.3 of the appendix.
To ensure a fair comparison, we compare our approach with existing methods by using the widely-used DIV2K dataset [53] as content reference and real LR images as degradation reference to create the paired training data for training various popular SR models. To maintain a rigorous and fair comparison, we have maintained consistency in all experimental settings and environments, with the exception of data generation. We utilize the public BasicSR  [^1] https://github.com/XPixelGroup/BasicSR for training Real-SR methods with 16 NVIDIA V100 GPUs.
Compared Data Generation Methods. We compare our methods with some state-of-the-art real-world data generation methods, including Hand-Crafted Physical-Based Degradation Models (BSRGAN [32] and Real-ESRGAN [33]) and learning-based degradation diffusion models proposed by [46], denoted as SynDiff for convenience.
Real-SR Models. To comprehensively validate the effectiveness of the generated data, we carefully select four classic and representative backbone architectures for evaluation, including representative CNN-based model RRDB [54], representative transformer-based model SwinIR [55] and HAT [56], representative lightweight model SwinIR-L [55]. To be fair, we conduct comparative evaluations under consistent experimental conditions and settings. We utilize L1Loss [56, 55] and perception loss [33, 57, 54] for training PNSR-oriented and Perception-oriented real SR models, respectively.
Metrics. For the PNSR-oriented Real SR model, we adopt PSNR [58] and SSIM [59] to quantitatively evaluate the performance. For the perception-oriented Real-SR model, we adopt LPIPS [60] and FID [61] to quantitatively evaluate the performance. DISTS [62] and CLIP-Score [63] are further introduced to evaluate the accuracy of generated data. Note that the higher the PSNR, SSIM, and CLIP-Score, the better, and the lower the LPIPS, FID, and DISTS, the better.
Evaluation. We utilize two public benchmarks to evaluate the performance of real-world image super-resolution methods, including the real-world dataset RealSR [28] and DRealSR [31] captured by digital single-lens reflex cameras (DSLRs). To further improve the diversity and quantity of real degradation scenarios, we have collected 891 pairs of data captured by smartphones for evaluation, denoted as SmartPhone.
@@
RealSR [28] DRealSR [31] SmartPhone
@@
SwinIR (Real-ESRGAN [33]) 24.395 0.7760 26.944 0.8308 27.395 0.8338
SwinIR (BSRGAN [32])) 25.852 0.7808 27.985 0.8308 28.049 0.8407
SwinIR (SynDiff [46]) 25.589 0.7687 28.301 0.8309 28.566 0.8453
@@
RRDB (Real-ESRGAN [33]) 24.579 0.7614 27.131 0.8193 27.841 0.8378
RRDB (BSRGAN [32])) 25.406 0.7685 27.523 0.8017 28.029 0.8278
RRDB (SynDiff [46]) 25.488 0.7691 28.078 0.8257 28.303 0.8426
@@
HAT (Real-ESRGAN [33]) 24.893 0.7726 27.339 0.8215 27.781 0.8336
HAT (BSRGAN [32])) 25.997 0.7816 28.135 0.8273 28.137 0.8369
HAT (SynDiff [46]) 25.790 0.7584 28.506 0.8286 28.508 0.8471
@@
SwinIR-L (Real-ESRGAN [33]) 24.367 0.7723 27.018 0.8244 27.581 0.8409
SwinIR-L (BSRGAN [32])) 25.651 0.7800 27.813 0.8301 28.118 0.8437
SwinIR-L (SynDiff [46]) 25.281 0.7516 28.170 0.8244 28.474 0.8487
@@
RealSR[28] DRealSR[31] SmartPhone
@@
SwinIR (Real-ESRGAN [33]) 0.3037 69.965 0.3219 39.175 0.4053 78.242
SwinIR (BSRGAN [32])) 0.2945 79.833 0.3023 38.541 0.3043 76.871
SwinIR (SynDiff [46]) 0.3835 103.179 0.3801 54.588 0.3129 83.485
@@
RRDB (Real-ESRGAN [33]) 0.3480 82.056 0.3551 39.310 0.3480 77.573
RRDB (BSRGAN [32])) 0.3041 77.412 0.3127 36.528 0.3381 78.812
RRDB (SynDiff [46]) 0.4004 98.798 0.4017 56.573 0.3511 88.431
@@
HAT (Real-ESRGAN [33]) 0.3066 79.209 0.3219 41.862 0.4022 87.950
HAT (BSRGAN [32])) 0.2852 80.192 0.2835 41.723 0.3049 81.247
HAT (SynDiff [46]) 0.3332 93.763 0.3465 50.808 0.3171 85.587
@@
SwinIR-L (Real-ESRGAN [33]) 0.3108 79.491 0.3234 42.986 0.4021 87.531
SwinIR-L (BSRGAN [32])) 0.3013 84.195 0.2978 43.246 0.3146 83.444

### v4

@@ 1 Introduction
Real-world image Super-Resolution (Real SR) is a fundamental problem in image processing, aiming to enhance the resolution and quality of images in real-world scenarios (Chen et al., 2022; Yu et al., 2024; Liu et al., 2023; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024). It has a wide range of applications across various fields, including photography (Chen et al., 2019) and medical imaging (Li et al., 2021), which enhance human visual perception and the robustness of vision systems (Haris et al., 2021; Noor et al., 2019; Gunturk et al., 2003; Chen et al., 2020a). However, traditional bicubic-interpolation-based Real SR methods have proven less effective in complex real-world scenarios due to the significant discrepancy between the bicubic pattern and real degradation (Chen et al., 2022; Liu et al., 2023; Cai et al., 2019; Wang et al., 2020; Chen et al., 2024). Consequently, substantial efforts have been directed towards developing methods for generating more realistic data to improve the generalization ability of Real SR models (Cai et al., 2019; Wei et al., 2020; Zhang et al., 2021; Wang et al., 2021b; Zhang et al., 2023a; Park et al., 2023; Li et al., 2022b; Xiao et al., 2020; Wolf et al., 2021; Luo et al., 2024; Sun & Chen, 2024; Hendrycks & Dietterich, 2019; Deng et al., 2023).
@@
To explore what kind of data contributes most to the SR model’s capability, we synthesize different sets of training data using (Elad & Feuer, 1997) and evaluate the performance of FSRCNN (Dong et al., 2016). As shown in Figure 1(a), the red triangles represent test samples of the target domain, while rectangular boxes indicate training sets with different blur kernels and noise levels. The results clearly demonstrate that the closer the training data distribution is to the test data, the better the model’s performance. This underscores the importance of designing a method that can adaptively generate accurate data for different target domains. Therefore, an ideal data generation system for Real SR should meet the following criteria: I) Large-scale, to satisfy the extensive data requirements for training deep learning models (Raicu et al., 2008; Maeda, 2020; Sehwag et al., 2022; Wu et al., 2023); II) Realistic, to enable Real SR models to accurately learn the characteristics of real-world degradation  (Cai et al., 2019; Wei et al., 2020; Zhang et al., 2021; Ji et al., 2020; Liu et al., 2022); and III) Adaptive, to flexibly generate data with arbitrary given degradation patterns, improving generalization in target domains (Liu et al., 2023; Zhang et al., 2023a; Chen et al., 2019; Lugmayr et al., 2019b; Mou et al., 2022).
@@
Existing data generation methods for Real SR (Cai et al., 2019; Wei et al., 2020; Zhang et al., 2021; Wang et al., 2021b; Maeda, 2020; Bulat et al., 2018; Yang et al., 2023; Lugmayr et al., 2020a; Yuan et al., 2018; Ignatov et al., 2017), as shown in Figure 1(b), can be broadly categorized into the following: a) Manual Collection via Focal Length Adjustment: This approach involves using digital single-lens reflex cameras (DSLRs) with varying focal lengths to capture images, followed by alignment (Cai et al., 2019; Wei et al., 2020). While it can produce realistic paired data, it is labor-intensive and often results in scene monotony and image misalignment, failing to meet the large-scale data requirement. b) Hand-Crafted Physical-Based Degradation Modeling: This method employs various degradation models (e.g., noise, blur, bicubic, JPEG) applied in single-order or higher-order combinations (Zhang et al., 2021; Wang et al., 2021b). Although efficient in generating large data quantities, the synthetic data often fails to accurately reflect the complex degradation patterns of real-world images, and its lack of adaptability to specific domains limits effectiveness. c) Learning-Based Methods: Techniques involving Generative Adversarial Networks (GANs) (Bulat et al., 2018) and diffusion models (Yang et al., 2023) are proposed to simulate realistic real-world degradation for low-resolution (LR) images. While these methods produce more realistic data compared to hand-crafted approaches, they often struggle to generalize to new and diverse domains, limiting their applicability in real-world scenarios. In summary, existing data generation methods face challenges in achieving both realism and adaptability: tailoring models to specific target domains may hinder their adaptability to new domains, and vice versa. Overcoming these challenges is crucial for advancing the field of Real SR.
Figure 1: (a) and (b) are SR performance on different train data and degradation distribution of different methods. (c) is the pipeline of our unsupervised data generation framework RealDGen.
@@
In this paper, we introduce a novel unsupervised learning framework Realistic Decoupled Data Generator (RealDGen) to meet the large-scale, realistic, and adaptive data generation criteria, as shown in Figure 1(c). RealDGen enhances degradation realism and content fidelity by separately modeling content and degradation through unsupervised learning and integrating them into a diffusion model to generate paired data. The training involves two steps: first, pre-training degradation and content extractors using contrastive and reconstruction learning to improve representation robustness; second, using these pre-trained extractors to condition the diffusion model with real LR degradation and content representations. To improve generalization to unknown LR distributions, the partial parameters of the extractors are fine-tuned. During data generation, unpaired HR and real LR images are used to extract content and degradation representations, which are then combined in the diffusion model. This process produces data that marries HR content with arbitrary LR degradation, improving adaptability to new domains. Extensive experiments show that RealDGen outperforms previous methods in generating realistic paired data and enhancing the performance of SR models in real-world scenarios.
@@
Figure 2: An overview of the training pipeline of our proposed RealDGen. We first train on the content and degradation extractors, then train Decoupled DDPM while fine-tuning the partial parameters of the extractors. RealDGen adaptively generates realistic LR images with arbitrarily given real LR images and unpaired HR images.
@@ 2.1 Content and Degradation Extractor Pre-training
To capture content and degradation representations, we propose dedicated degradation and content extractors, denoted as $E_{deg}$ and $E_{cont}$, respectively. We employ reconstruction learning for training $E_{cont}$, as shown in the left of Figure 2. Specifically, for a given high-resolution (HR) image $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, we degrade it to a low-resolution (LR) counterpart $\mathcal{X}_{lr}\in\mathbb{R}^{C\times h\times w}$ by Real-ESRGAN (Wang et al., 2021b) degradation model $\mathcal{D}$ with diverse synthetic degradations. Subsequently, $E_{cont}$ is engaged to extract the content representation $F_{{cont}}$ from $\mathcal{X}_{lr}$. Thereafter, a HR reconstruction network $Rec_{hr}$ is harnessed to reconstruct the HR image $\hat{\mathcal{X}}\in\mathbb{R}^{C\times H\times W}$ from $F_{{cont}}$, as follows:
@@
$\mathcal{L}_{{rh}}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{\mathcal{X}}_{i}-\mathcal{X}_{i}\right)^{2}\,$ (2)
where $N$ denotes the batch size, which we empirically set to 64. After using LR-HR paired training, $E_{cont}$ is able to learn a robust content representation under diverse degradation and real scenarios.
Considering the variability of degradation in diverse scenarios and imaging devices in the real world, we advocate for contrastive learning (Chen et al., 2020b; Hermans et al., 2017) to curate positive and negative samples for training $E_{deg}$, which guarantees the uniqueness of the degradation representations. Specifically, for a HR image $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, we generate an LR image $\mathcal{X}_{{lr}}\in\mathbb{R}^{C\times h\times w}$ by $\mathcal{D}$ with parameter $\theta$ as the anchor $\mathcal{X}_{{anc}}$. We further obtain a set of negative samples $\mathcal{X}_{{neg}_{i}}\in\mathbb{R}^{C\times h\times w}$ by applying $\mathcal{D}$ with different parameters $\theta_{i}$ to $\mathcal{X}$, and a set of positive samples $\mathcal{X}_{{pos}_{i}}\in\mathbb{R}^{C\times h\times w}$ by applying the degradation $\mathcal{D}$ with same parameter $\theta$ to different HR images $\mathcal{X}^{\prime}_{i}\in\mathbb{R}^{C\times H\times W}$, as follows:
$\begin{array}[]{c}\mathcal{X}_{{neg}}=\{\mathcal{D}\left(\mathcal{X},\theta_{1}\right),\mathcal{D}\left(\mathcal{X},\theta_{2}\right),\ldots,\mathcal{D}\left(\mathcal{X},\theta_{n}\right)\},\\ \mathcal{X}_{{pos}}=\{\mathcal{D}\left(\mathcal{X}^{\prime}_{1},\theta\right),\mathcal{D}\left(\mathcal{X}^{\prime}_{2},\theta\right),\ldots,\mathcal{D}\left(\mathcal{X}^{\prime}_{n},\theta\right))\}\,.\end{array}$ (3)
@@
$\mathcal{L}_{{cl}}=\sum_{i=1}^{n}\max\left(0,d\left(E_{deg}\left(\mathcal{X}_{{anc}}\right),E_{deg}\left(\mathcal{X}_{{pos}_{i}}\right)\right)-d\left(E_{deg}\left(\mathcal{X}_{{anc}}\right),E_{deg}\left(\mathcal{X}_{{neg}_{i}}\right)\right)+{margin}\right)\,.$ (4)
where $d$ symbolizes the L2 distance, $n$ is the number of samples, and we empirically set $n$ and $margin$ to $3$ and $0.01$, respectively. Furthermore, to drive $E_{deg}$ to learn the complete degradation representations, we utilize the reconstruction learning strategy as aforementioned for supervising $E_{deg}$. Specifically, we utilize the pre-trained $E_{cont}$ to learn the content representation of HR images $\mathcal{X}_{{hr}}$ and $E_{deg}$ to learn the degradation representations of $\mathcal{X}_{{lr}}$. We employ a low-resolution reconstruction network, $Rec_{lr}$, to combine these representations and reconstruct the LR image, $\hat{\mathcal{X}}_{{lr}}$, as follows:
@@
After training with well-designed learning strategies, our extractors effectively capture robust content and degradation representations. More details and analysis are provided in Appendix A.1 and A.2.
@@
5:  $F_{{cont}}=E_{cont}\left(\mathbf{x}_{lr}\right)$
6:  ${F_{deg}}=E_{deg}\left(\mathbf{x}_{lr}\right)$
7:  $\mathbf{c}=\mathcal{M}\left({F_{cont}},{F_{deg}}\right)$
@@ 2.2 Decoupled DDPM Training and Extractor Fine-tuning
We introduce a content-degradation Decoupled Diffusion Probabilistic Model (Decoupled DDPM) to generate real LR images. In detail, given real LR images ${\mathcal{X}}_{{lr}}$ from real-world $q$ encompassing various degradations, we extract their robust content representation $F_{cont}$ and degradation representation $F_{deg}$ by pre-trained $E_{cont}$ and $E_{deg}$, respectively. To enhance the generalization of $E_{deg}$ and $E_{cont}$ on unseen real distributions, partial parameters are fine-tuned, as shown in Figure 2. A modulation block $\mathcal{M}$ is introduced to adequately incorporate degradation representation into the content, formulated as:
$\mathbf{c}=\mathcal{M}\left(E_{deg}\left(\mathbf{x}_{lr}\right),E_{cont}\left(\mathbf{x}_{lr}\right)\right)\,.$ (7)
Then, this fused image representation is utilized as a condition $\mathbf{c}$ for controlling our Decouple DDPM to generate LR images. To make it clear, we illustrate the detailed training procedure of the Decouple DDPM, as shown in Algorithm 1. More details of fine-tuning and analysis of $E_{cont}$ and $E_{deg}$ are provided in Sections A.3A.4 and A.5 of the appendix.
@@ 2.3 Data Generation
We propose a novel strategy to generate realistic LR images using unpaired LR and HR images by decoupling content and degradation. First, we extract the degradation representation from a real-world LR image and the content representation from an HR image. These representations are combined in the modulation module $\mathcal{M}$ to serve as the condition $\mathbf{c}$ for the diffusion model to generate LR images. The generated LR images retain the content of the HR image and the degradation of the real LR image, as shown on the right of Figure 1. To enhance fidelity, following (Meng et al., 2021), we denoise from an initial LR image $\mathbf{x}_{t}$ with $t$ steps of noise rather than from pure Gaussian noise. This initial LR image is degraded by ${\mathcal{D}}$. In Section 3.4, we analyze the step number $T$. Details of our data generation pipeline are in Algorithm 2.
Although using content and degradation conditions improves the controllability and fidelity of Decoupled DDPM, the inherent stochasticity of the diffusion model (Ho et al., 2020; Rombach et al., 2022) can still introduce tiny artifacts and content distortion. To mitigate this, we propose a filtering mechanism. For each generated LR image, we re-extract content and degradation representations, then calculate the degradation error with the real LR image and the content error with the HR image. By selecting samples with the smallest errors, we reduce diffusion stochasticity and produce higher-fidelity LR images. More details and analysis are presented in Appendix A.3A.5 and A.6.
@@ 3.1 Experiments Settings
Training details. We collect about 152,000 real low-resolution images from both public datasets (Wei et al., 2020; Cai et al., 2019; Ignatov et al., 2017) and those captured using smartphones to train RealDGen. Extractors and Decouple DDPM are trained with learning rate $1\times 10^{-4}$ and batch size $16$ on 16 NVIDIA V100 GPUs. More details of the training setting are presented in Section A.1 and A.3 of the appendix. To ensure a fair comparison, we compare our approach with existing methods by using the widely-used DIV2K dataset (Agustsson & Timofte, 2017) as HR images and real LR images as degradation references to create the paired training data for training various popular SR models. To maintain a rigorous and fair comparison, we maintain consistency in all experimental settings and environments, with the exception of data generation. We utilize the public BasicSR for training Real-SR methods with 16 NVIDIA V100 GPUs.
Compared data generation methods. We compare our methods with some state-of-the-art real-world data generation methods, including Hand-Crafted Physical-Based Degradation Models (BSRGAN (Zhang et al., 2021) and Real-ESRGAN (Wang et al., 2021b)) and learning-based degradation diffusion models proposed by (Yang et al., 2023), denoted as SynDiff for convenience.
Real-SR models. To comprehensively validate the effectiveness of the generated data, we select five classic and representative backbone architectures for evaluation, including CNN-based model RRDB (Wang et al., 2018), transformer-based model SwinIR (Liang et al., 2021) and HAT (Chen et al., 2023b), diffusion-based model ResShift (Yue et al., 2024) and lightweight model SwinIR-L (Liang et al., 2021). To be fair, we conduct comparative evaluations under consistent experimental conditions and settings. We utilize L1Loss (Chen et al., 2023b; Liang et al., 2021) and perception GAN loss (Wang et al., 2021b; Johnson et al., 2016; Wang et al., 2018) for training PNSR-oriented and Perception-oriented Real SR models, respectively.
Metrics. For the PNSR-oriented Real SR model, we adopt PSNR (Huynh-Thu & Ghanbari, 2008) and SSIM (Wang et al., 2004) to quantitatively evaluate the performance. For the perception-oriented Real-SR model, we adopt LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) to quantitatively evaluate the performance. DISTS (Ding et al., 2020) and CLIP-Score (Radford et al., 2021) are further introduced to evaluate the accuracy of generated data. Note that the higher the PSNR, SSIM, and CLIP-Score, the better, and the lower the LPIPS, FID, and DISTS, the better.
@@
RealSR DRealSR SmartPhone
@@
SwinIR (Real-ESRGAN) 24.395 0.7760 26.944 0.8308 27.395 0.8338
SwinIR (BSRGAN) 25.852 0.7808 27.985 0.8308 28.049 0.8407
SwinIR (SynDiff) 25.589 0.7687 28.301 0.8309 28.566 0.8453
@@
RRDB (Real-ESRGAN) 24.579 0.7614 27.131 0.8193 27.841 0.8378
RRDB (BSRGAN) 25.406 0.7685 27.523 0.8017 28.029 0.8278
RRDB (SynDiff) 25.488 0.7691 28.078 0.8257 28.303 0.8426
@@
HAT (Real-ESRGAN) 24.893 0.7726 27.339 0.8215 27.781 0.8336
HAT (BSRGAN) 25.997 0.7816 28.135 0.8273 28.137 0.8369
HAT (SynDiff) 25.790 0.7584 28.506 0.8286 28.508 0.8471
@@
SwinIR-L (Real-ESRGAN) 24.367 0.7723 27.018 0.8244 27.581 0.8409
SwinIR-L (BSRGAN) 25.651 0.7800 27.813 0.8301 28.118 0.8437
SwinIR-L (SynDiff) 25.281 0.7516 28.170 0.8244 28.474 0.8487
@@
RealSR DRealSR SmartPhone
@@
SwinIR (Real-ESRGAN) 0.3037 69.965 0.3219 39.175 0.4053 78.242
SwinIR (BSRGAN) 0.2945 79.833 0.3023 38.541 0.3043 76.871
SwinIR (SynDiff) 0.3835 103.179 0.3801 54.588 0.3129 83.485
@@
RRDB (Real-ESRGAN) 0.3480 82.056 0.3551 39.310 0.3480 77.573
RRDB (BSRGAN) 0.3041 77.412 0.3127 36.528 0.3381 78.812
RRDB (SynDiff) 0.4004 98.798 0.4017 56.573 0.3511 88.431
@@
HAT (Real-ESRGAN) 0.3066 79.209 0.3219 41.862 0.4022 87.950
HAT (BSRGAN) 0.2852 80.192 0.2835 41.723 0.3049 81.247
HAT (SynDiff) 0.3332 93.763 0.3465 50.808 0.3171 85.587
@@
SwinIR-L (Real-ESRGAN) 0.3108 79.491 0.3234 42.986 0.4021 87.531
SwinIR-L (BSRGAN) 0.3013 84.195 0.2978 43.246 0.3146 83.444